OpenStudy (anonymous):

In a study of 250 adults, the mean heart rate was 70 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute. What is the 99% confidence interval for the mean beats per minute? (5 points) 68.9 - 76.3 70 - 72 61.2 - 72.8 68 - 72

OpenStudy (anonymous):

someone please help!

OpenStudy (kropot72):

A 99% confidence interval for the mean is: \[\large (\bar{x}-2.576\frac{\sigma}{\sqrt{n}},\ \bar{x}+2.676\frac{\sigma}{\sqrt{n}})\] Now you just need to plug the given values into the formula.

OpenStudy (anonymous):

So \[\sigma\] would be 12 but which would be x and which would be n? I know its between 250 and 70 but I dont know which one is which? @kropot72 @tkhunny

OpenStudy (kropot72):

n is the sample size. In this case n = 250.

OpenStudy (tkhunny):

\(\sigma = 12\)

OpenStudy (kropot72):

x-bar is 70.

OpenStudy (kropot72):

A correct to my posted formula. the z-critical value is 2.567 in both terms.

OpenStudy (anonymous):

So it would be 68-72 right?

OpenStudy (kropot72):

Correct! Good work.

OpenStudy (anonymous):

Thank you so much!! I understand it now!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (tkhunny):

kropot72 was paying better attention. Leave ½% in each tail. :-)